A high school computer club needs to raise $1,200.00 divided equally amongst the members. When 8 new members join, the individual amount is reduced by $7.50. How many members are in the club?
Here is what I have done so far:
X is the original number of members and X + 8 is the new number of members.
The total amount to raise never changes. 1,200/X is the original amount each member would need to contribute and 1,200/X + 8 is the new amount.
I am not sure how to set this up and solve to find the number of members. Your help would be much appreciated. Thanks

Answer by checkley75(3666)

1200/X-1200/(X+8)=7.50
[1200(X+8)-1200X]/X(X+8)=7.50
(1200X+9600-1200X)/X^2+8X=7.50
7.5(X^2+8X)=9600
7.5X^2+60X-9600=0
(X-32)(7.5X+300)=0
X-32=0
X=32 ORIGINAL MEMBERS.
PROOF
1200/32-1200/40=7.5
37.5-30=7.5
7.5=7.5